The mass of a substance varies directly as the volume of the substance. If the mass of 7 liters of the substance is 42 kilograms, what will be the volume in liters of 63 kilograms of the substance?
Found 2 solutions by math_helper, josgarithmetic:
Answer by math_helper(2461)


You can calculate the density (mass per unit volume) easily enough:
42kg/7l = 6kg/l
Now you can use the density to compute volume of a given mass, by dividing:
V = 63kg / (6kg/l) = 10.5l

---
For this problem, a proportion would also work:
7/42 = x/63 --> x = 63*7/42 = 10.5

Answer by josgarithmetic(39617)

Density is being taken as a constant.